How do you take a multiple choice test?
Multiple choice tests can be quite easy or quite hard. Here are some tips for learning how to take them:
Give examples of two events that are independent?
the result of throwing a dice and the result of taking a card from a deck at the same time
A 6-sided dice is rolledwhat is the probability of rolling a 3 or 4?
The probability of rolling a three or a four (meaning one roll, and either number would be counted a "success") is 1/3.
This is because the probability of rolling a three is 1/6, the probability of rolling a four is 1/6, and if you are good with either possibility, then the combined probability is 1/6+1/6 = 1/3.
What are the possible outcomes when you roll a number cube?
1,2,3,4,5,6 are the possible outcomes. There's a 1/6 chance for each.
What is the probability that 16 rolls of a fair die will show 4 ones?
evens says 4 * 6 = 24 rolls required for 4 shows
so 4 shows in 16 = 16 : 24 = 5 to 8 against
A z-score is a means to compare rank from 2 different sets of data by converting the individual scores into a standard z-score. The formula to convert a value, X, to a z-score compute the following: find the difference of X and the mean of the date, then divide the result by the standard deviation of the data.
What is the x deck card in uno attack?
It's called a "Discard all" card.
- When played, you must discard all cards of the color of the discard all card
(E.g: If the top card on the discard pile is a blue 4, the player may play a blue discard all card and discard all their blue cards.)
100%
What probability cannot happen?
By definition, the probability of an event is always some number between 0 and 1. Sometimes we think of these as percents. For example, if something has a probability of 25%, the probability is simply 25/100, or .25.
By this definition, nothing can have probability 3 or 1.2, or any negative number whatsoever. Only probabilities between 0 and 1 (inclusive) are allowed.
What is the probablity that the sequence will be heads-heads-heads-heads if you toss a coin 4 times?
Each independent trial has a 1/2 probability that a heads will result. So for a sequence of 4 heads you have 1/2 x 1/2 x 1/2 x 1/2 = 1/16
What is the event of one outcome happening and that outcome not happening?
It is an impossible event - which has probability zero.
What is the important law of binomial distribution?
1 - 2 answers
2 - probabilty of success remains constant
3 - outcomes of trials are independent
4 - random variable of interest is the number of successes in a fixed number of of trials
What is 3.5 billion as a percentage?
Your question is actually meaningless without context. But allow me to propose one. Let us say that the question is, 3.5 billion is what percentage of the current total human population of nearly 7 billion people. The answer then is nearly 50%.
How are experimental probability and theoretical probability related?
As the number of times that the experiment is conducted increases, the experimental probability will near the theoretical probability - unless there is a problem with the theoretical model.
What are the chances of injury to 'vas deferens' in open hernia surgery?
Unless the vas deference has herniated also, the chances are almost zero.
The probability of rolling a multiple of five on a standard die is 1 in 6, or about 0.1667.
The probability of rolling a 10, 15, or higher is zero, because the question implied only one die.
What important question must you answer before computing an and probability?
Whether or not the "and" events are independent.
How do you make a cumulative frequency table?
Given a frequency table,
An alternative definition is that the cumulative frequency for any value is the sum of all the frequencies less than or equal to that value.
If you pull a card from a deck of cards what is the propability of getting a king or a heart?
0.2884615
How likely is it to have two miscarriage in a roll?
it is very possible to have two or more misscariages one after the other, i had a misscariage in June at 9 weeks pregnant and am now 7 weeks pregnant and the same thing is happening again, it is now November! hope that has helped good luck xx